# Fourier dimension at $C^{1+\alpha}$ regularity

### v2

@@ 2.2. Moment bounds from the tail condition
where $\tau:I\to\mathbb{R}$ is the geometric potential $\tau(x)=-\ln|T^{\prime}(x)|$, $x\in I$. Proposition 2.1 quantifies deviation from this scale of contraction using properties of the Gurevich pressure $P(\phi+t(\tau-\lambda))$, which is defined for a potential $\psi$ as
$P(\psi):=\lim_{n\to\infty}\frac{1}{n}\log\sum_{\mathbf{a}\in\Sigma^{n}}\sup_{x\in I_{b(\mathbf{a})}}e^{S_{n}\psi(g_{\mathbf{a}}(x))},$
where $S_{n}\psi(x)=\sum_{j=0}^{n-1}\psi(T^{k}(x))$. Notice that since $\mathcal{L}_{\phi}^{n}(1)=1$, we find that our normalized potential $\phi$ satify $P(\phi)=0$. It is also known that $\frac{d}{dt}P(\phi+t\tau)_{|t=0}=\int\tau d\mu$. The differentiability properties of the pressure allows us to bound the spreading of the random variable $e^{\lambda n}|g_{\mathbf{a}}^{\prime}(x)|$ in the following way:
@@
The proof of Proposition 2.1 can be deduced from the proof leading to large deviation principle, e.g. [25] and the references there-in. However, as our result is stated quantitatively in terms of moments, we included a direct proof using Sarig’s results from [43] for completeness.
@@
Let $\gamma_{0}\leq t<0$. Then if write $p_{a}:=\mu(I_{a})$ and recalling that $\kappa_{-}(a)\leq|g_{a}^{\prime}(x)|\leq\kappa_{+}$ for all $x\in I_{b}$, where we have supressed the prefix notation of $b$ for admissibility, then by the light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail property and using the Gibbs constant $C_{\phi}$ we have
@@
Since $\tau$ is $\alpha$-Hölder by the bounded distortion property and as the pressure $P(\varphi+t(\tau-\lambda))$ is finite for all $|t|\leq\gamma_{0}$, we know by [43, Theorem 4] that there exists $0<t_{\gamma_{0}}^{\prime}<\gamma_{0}$ such that the function $f:t\in(-t_{\gamma_{0}}^{\prime},t_{\gamma_{0}}^{\prime})\mapsto P(\varphi+t(\tau-\lambda))$ is real analytic in $(-t_{\gamma_{0}}^{\prime},t_{\gamma_{0}}^{\prime})$ and satisfies that $f(0)=P(\varphi)=0$ and $f^{\prime}(0)=\int_{I}(\tau-\lambda)d\mu=0$. Taylor estimate near 0 then gives some $t_{\gamma_{0}}^{\prime\prime}\in(0,1)$ such that $|f(t)|\leq\tfrac{3}{4}f^{\prime\prime}(0)t^{2}$ for $|t|\leq t_{\gamma_{0}}^{\prime\prime}$. Thus choosing $0<t_{\gamma_{0}}<t_{\gamma_{0}}^{\prime\prime}$ small enough, it follows that, for all $|t|\leq t_{\gamma_{0}}$, there exists $n_{0}(t)\geq 1$ such that
@@ 2.3. Upper regularity of the Gibbs measure from the tail
The moment bound deduced in Proposition 2.1 allows us to prove the following Frostman condition or upper regularity bound for the measure $\mu$ allowing us to disregard $\mu$ concentrating near points at small scales. We note that if one would relax the light tail assumption to heavier tail assumptions, the upper regularity of $\mu_{\phi}$ is likely to fail.
@@
There exists $C_{reg}\geq 1$, $s_{\mu}\in(0,1)$ such that:
$\forall x\in\mathbb{R},\ \forall\sigma\in(0,1),\ \mu([x-\sigma,x+\sigma])\leq C_{reg}\sigma^{s_{\mu}}.$
@@
First of all, $\sum_{\mathbf{a}\in\Sigma^{n}}\mu(I_{\mathbf{a}})=\mu(I)$, so that $(I_{\mathbf{a}})_{\mathbf{a}}$ is a partition of $I$ up to a zero-mesure set. Now consider some interval $J\subset I$. Define $\mathcal{W}_{n}(J):=\{\mathbf{a}\in\Sigma^{n},\ I_{\mathbf{a}}\cap J\neq 0\}$. We have, for all $n\geq 1$:
$J\subset\bigcup_{\mathbf{a}\in\mathcal{W}_{n}(J)}I_{\mathbf{a}}.$
Hence:
$\mu(J)\leq\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)}\mu(I_{\mathbf{a}})\lesssim\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)}w_{\mathbf{a}}(x_{b(\mathbf{a})}).$
for any choice of reference points $x_{a}\in I_{a}$. Define further: $E_{n}:=\{\mathbf{a}\in\Sigma^{n},\ |g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|\geq e^{-2\lambda n}\}.$ Proposition 2.1 then yields, for $t$ small enough and $n$ large enough (depending on $t$):
$\sum_{\mathbf{a}\in\Sigma^{n}\setminus E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})\leq\sum_{\mathbf{a}\in\Sigma^{n}\setminus E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})e^{-2\lambda nt}|g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|^{-t}=e^{-\lambda nt}\sum_{\mathbf{a}\in\Sigma^{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})|e^{\lambda n}g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|^{-t}$
$\leq 2C_{\phi}e^{-\lambda nt}e^{\varepsilon_{\gamma_{0}}nt^{2}}\leq 2C_{\phi}e^{-n\tilde{\varepsilon}}$,
for some $\tilde{\varepsilon}>0$. We can then write:
$\mu(J)\lesssim\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)}w_{\mathbf{a}}(x_{b(\mathbf{a})})\lesssim\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})+e^{-\tilde{\varepsilon}n}.$
Our second step is to bound $w_{\mathbf{a}}$ by a power of $|g_{\mathbf{a}}^{\prime}|$. First of all, recall that our moment condition is $\sum_{a\in\Sigma}\sup_{x}w_{ab}(x)|g_{ab}^{\prime}(x)|^{-\gamma}<\infty$ for some $\gamma>0$. It follows that there exists a constant $C$ such that:
$\forall ab\in\Sigma^{2},\forall x\in I_{b},\ \ w_{ab}(x)\leq C|g_{ab}^{\prime}(x)|^{\gamma}.$
Now, since $\sum_{ab\in\Sigma}|I_{ab}|<\infty$, we know that $g_{ab}^{\prime}(x)$ goes to zero when the word $ab$ goes far in the set $\Sigma^{2}$. In particular, there exists a finite set $\mathcal{B}\subset\Sigma^{2}$ such that:
$\forall ab\in\Sigma^{2}\setminus\mathcal{B},\ w_{ab}(x)\leq|g_{ab}^{\prime}(x)|^{\gamma/2}.$
Then, for every $ab\in\mathcal{B}$, we can write, since $\varphi\leq-\varepsilon_{\varphi}<0$:
$w_{ab}(x)=\exp(\varphi(g_{ab}(x)))\leq\exp(-\varepsilon_{\varphi})\leq\exp\Big(-\frac{\varepsilon_{\varphi}}{\max_{cd\in\mathcal{B}}|\ln g_{cd}^{\prime}|_{\infty}}\big|\ln|g_{ab}^{\prime}(x)|\big|\Big)=|g_{ab}^{\prime}(x)|^{\widehat{\varepsilon}}$,
for some $\widehat{\varepsilon}\in(0,1)$. We then find, reducing $\widehat{\varepsilon}$ if necessary:
$\forall ab\in\Sigma^{2},\forall x\in I_{b},\ w_{ab}(x)\leq|g_{ab}^{\prime}(x)|^{-\widehat{\varepsilon}}$.
We can then bound similar quantities for some word $\mathbf{a}$. We find, for any $\mathbf{a}\in\Sigma^{n}$, and for any $x\in I_{b(\mathbf{a})}$:
$w_{\mathbf{a}}(x)=w_{a_{1}a_{2}}(g_{a_{2}\dots a_{n}}(x))\dots w_{a_{n-1}a_{n}}(x)\leq|g_{a_{1}a_{2}}^{\prime}(g_{a_{2}\dots a_{n}}(x))^{\widehat{\varepsilon}}\dots g_{a_{n-1}a_{n}}^{\prime}(x)^{\widehat{\varepsilon}}|=|g_{\mathbf{a}}^{\prime}(x)|^{\widehat{\varepsilon}}$.
We then inject this bound into the previous sum to find, using Hölder’s inequality:
$\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})\leq\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}|g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|^{\widehat{\varepsilon}}\leq\Big(\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}|g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|\Big)^{\widehat{\varepsilon}}\cdot\#(\mathcal{W}_{n}(J)\cap E_{n})^{1-\widehat{\varepsilon}}$.
Now notice the following. Since we are working with a partition of intervals, there can only exists at most $2$ words $\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}$ such that $I_{\mathbf{a}}\nsubseteq J$. So we can write
$\mathcal{W}_{n}(J)\cap E_{n}=X_{n}(J)\cup Y_{n}(J),$
where $\bigcup_{\mathbf{a}\in X_{n}(J)}I_{\mathbf{a}}\subset J$, and where $Y_{n}(J)\leq 2$. We can then bound the cardinal of $X_{n}(J)$, as follows:
$X_{n}(J)\leq e^{\lambda n/2}\sum_{\mathbf{a}\in X_{n}(J)}|I_{\mathbf{a}}|\leq|J|e^{\lambda n/2}.$
We then relate $n$ with $|J|$ by asking for this quantity to be bounded:
$n:=-\frac{2}{\lambda}\ln(|J|).$
In this case, $\mathcal{W}_{n}(J)\cap E_{n}\lesssim 1$. We finally conclude the proof in the following way, denoting
$\kappa=\sup_{a\in\Sigma}\|g_{a}^{\prime}\|_{\infty}\in(0,1)$:
$\mu(J)\lesssim\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}w_{\mathbf{a}}(x_{b(\mathbf{a})})+e^{-\tilde{\varepsilon}n}\lesssim\Big(\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}|g_{\mathbf{a}}^{\prime}(x_{b(\mathbf{a})})|\Big)^{\widehat{\varepsilon}}+e^{-\tilde{\varepsilon}n}$
$\lesssim\Big(\sum_{\mathbf{a}\in\mathcal{W}_{n}(J)\cap E_{n}}|I_{\mathbf{a}}|\Big)^{\widehat{\varepsilon}}+e^{-\tilde{\varepsilon}n}=\Big(2\kappa^{n}+\sum_{\mathbf{a}\in X_{n}(J)}|I_{\mathbf{a}}|\Big)^{\widehat{\varepsilon}}+e^{-\tilde{\varepsilon}n}$
$\lesssim\kappa^{\widehat{\varepsilon}n}+|J|^{\widehat{\varepsilon}}+e^{-\tilde{\varepsilon}n}\lesssim|J|^{s_{\mu}},$
for some $s_{\mu}>0$ small enough. ∎

### v3

@@ 2.2. Moment bounds from the tail condition
where $\tau:I\to\mathbb{R}$ is the geometric potential $\tau(g_{\mathbf{a}}(x))=\ln|g_{\mathbf{a}}^{\prime}(x)|$, $x\in I$, $\mathbf{a}\to x$. Proposition 2.1 quantifies deviation from this scale of contraction using properties of the Gurevich pressure $P(\phi+t(\tau-\lambda))$, which is defined for a potential $\psi$ as
$P(\psi):=\lim_{n\to\infty}\frac{1}{n}\log\sum_{\mathbf{a}\in\Sigma^{n}}\sup_{x\in I_{b(\mathbf{a})}}e^{S_{n}\psi(g_{\mathbf{a}}(x))}.$
Notice that since $\mathcal{L}_{\phi}^{n}(1)=1$, we find that our normalized potential $\phi$ satify $P(\phi)=0$. It is also known that $\frac{d}{dt}P(\phi+t\tau)_{|t=0}=\int\tau d\mu$. The differentiability properties of the pressure allows us to bound the spreading of the random variable $e^{\lambda n}|g_{\mathbf{a}}^{\prime}(x)|$ in the following way:
@@
The proof of Proposition 2.1 can be deduced from the proof leading to large deviation principle, e.g. [38] and the references there-in. However, as our result is stated quantitatively in terms of moments, we included a direct proof using Sarig’s results from [69] for the countable Markov shift for completeness.
@@
Let $\gamma_{0}\leq t<0$. Then if write $p_{a}:=w_{a}(x_{b}(a))$ and recalling that $\kappa_{-}(a)\leq|g_{a}^{\prime}(x)|\leq\kappa_{+}$ for all $x\in I_{b}$, where we have supressed the prefix notation of $b$ for admissibility, then by the light $(C_{\phi,\gamma_{0}},\gamma_{0})$-tail property and using the Gibbs constant $C_{\phi}$ we have
@@
Since $\tau$ is $\alpha$-Hölder by the bounded distortion property and as the pressure $P(\varphi+t(\tau-\lambda))$ is finite for all $|t|\leq\gamma_{0}$, we know by [69, Theorem 4] that there exists $0<t_{\gamma_{0}}^{\prime}<\gamma_{0}$ such that the function $f:t\in(-t_{\gamma_{0}}^{\prime},t_{\gamma_{0}}^{\prime})\mapsto P(\varphi+t(\tau-\lambda))$ is real analytic in $(-t_{\gamma_{0}}^{\prime},t_{\gamma_{0}}^{\prime})$ and satisfies that $f(0)=P(\varphi)=0$ and $f^{\prime}(0)=\int_{I}(\tau-\lambda)d\mu=0$. Taylor estimate near 0 then gives some $t_{\gamma_{0}}^{\prime\prime}\in(0,1)$ such that $|f(t)|\leq\tfrac{3}{4}f^{\prime\prime}(0)t^{2}$ for $|t|\leq t_{\gamma_{0}}^{\prime\prime}$. Thus choosing $0<t_{\gamma_{0}}<t_{\gamma_{0}}^{\prime\prime}$ small enough, it follows that, for all $|t|\leq t_{\gamma_{0}}$, there exists $n_{0}(t)\geq 1$ such that
@@ 2.3. Upper regularity of the Gibbs measure from the tail
We now prove the Frostman condition under the hypothesis that has the light-tail condition. I We note that if one would relax the light tail assumption to heavier tail assumptions, the upper regularity of $\mu_{\phi}$ is likely to fail. The result here is similar to that for overlapping self-similar measures by Feng and Lau [33] but generalised to the conformal setting and with infinite branches.
@@
Suppose that the attractor to the IFS $F$ is not a single point. Then assume one of the following:
  • the IFS has no overlaps, or
  • the IFS has overlaps but the potential $\phi:\Sigma^{*}\rightarrow\mathbb{R}$ is locally constant in the sense that $\omega_{a}(x)=p_{a}$, with $\sum_{a}p_{a}=1$
Then exists $C_{reg}\geq 1$, $s_{\mu}\in(0,1)$ such that:
$\forall I\text{ interval},\ \mu(I)\leq C_{reg}|I|^{s_{\mu}}.$
We note that all the examples we apply Theorem 1.1 satisfy either of the assumptions (e.g. the induced overlapping infinite IFS associated to the Lyons IFS in Theorem 1.3 satisfies the locally constancy).
To prove Proposition 2.2, we first need a folklore result about the absence of atoms of equilibrium states under the assumptions in Proposition 2.2:
Lemma 2.3.
Suppose that $F$ is not a single point.
  • If the IFS have no overlaps, then equilibrium states have no atoms.
  • If the IFS have overlaps, and if the potential $\phi:\Sigma^{*}\rightarrow\mathbb{R}$ is locally constant in the sense that $\omega_{a}(x)=p_{a}$, with $\sum_{a}p_{a}=1$, then $\mu_{\phi}$ have no atoms.
@@
Suppose that $\mu_{\phi}$ have atoms. Consider a maximal atom $x_{0}\in F$ of mass $M>0$. Then, for all $n$:
$M=\mu(x_{0})=\int_{F}\mathbf{1}_{x_{0}}d\mu(x)=\int_{F}\mathcal{L}_{\varphi}^{n}(\mathbf{1}_{x_{0}})d\mu=\sum_{\mathbf{a}\in\Sigma^{n}}\int_{F}\omega_{\mathbf{a}}(y)\mathbf{1}_{x_{0}}(g_{\mathbf{a}}(y))d\mu(y)$
$\underset{I_{\mathbf{a}}\ni x_{0}}{\sum_{\mathbf{a}\in\Sigma^{n}}}\int_{F}\omega_{\mathbf{a}}(y)\mathbf{1}_{g_{\mathbf{a}}^{-1}x_{0}}(y)d\mu(y)=\underset{I_{\mathbf{a}}\ni x_{0}}{\sum_{\mathbf{a}\in\Sigma^{n}}}\omega_{\mathbf{a}}(g_{\mathbf{a}}(x_{0}))\mu({g_{\mathbf{a}}^{-1}x_{0}})\leq M\underset{I_{\mathbf{a}}\ni x_{0}}{\sum_{\mathbf{a}\in\Sigma^{n}}}\omega_{\mathbf{a}}(g_{\mathbf{a}}(x_{0}))\mu({g_{\mathbf{a}}^{-1}x_{0}}).$
Hence, we get the key bound:
$1\leq\underset{I_{\mathbf{a}}\ni x_{0}}{\sum_{\mathbf{a}\in\Sigma^{n}}}\omega_{\mathbf{a}}(g_{\mathbf{a}}(x_{0})).$
Notice that $\omega_{\mathbf{a}}(x)<e^{-\varepsilon|a|}$ for some $\varepsilon$, hence this inequality can not hold if the IFS doesn’t have overlaps, thus proving our first point. In the case where $F$ have overlaps, if we further suppose that $\omega_{\mathbf{a}}(y)$ are constants, we find:
$1\leq\underset{I_{\mathbf{a}}\ni x_{0}}{\sum_{\mathbf{a}\in\Sigma^{n}}}\omega_{\mathbf{a}}\leq{\sum_{\mathbf{a}\in\Sigma^{n}}}\omega_{\mathbf{a}}=1$
and hence $\forall n,\forall\mathbf{a}\in\Sigma^{n},\ I_{\mathbf{a}}\ni x_{0}$. The fact that $F=\bigcap_{n}\bigcup_{\mathbf{a}\in\Sigma^{n}}\overline{I_{\mathbf{a}}}$ and that $|I_{\mathbf{a}}|\leq e^{-\varepsilon n}$ for some $\varepsilon>0$ implies that $F=\{x_{0}\}$. ∎
We now apply Lemma 2.3 to get upper regularity via an induction-by-scales strategy inspired by Feng and Lau [33]. Firstly we need the following definition and a technical lemma.
Definition 2.4.
Define
$\text{Reg}(F):=\{h:K\rightarrow\mathbb{R}_{+}^{*}\ |\ \forall x,y\in F,\forall\mathbf{b}\in\Sigma^{*},\omega_{\mathbf{b}}(x)h(g_{\mathbf{b}}(x))\leq C_{\phi}\omega_{\mathbf{b}}(y)h(g_{\mathbf{b}}(y))\},$
where $C_{\phi}$ is the Gibbs constant. Notice that $1\in\text{Reg}(F)$ by Gibbs estimates.
Lemma 2.5.
There exists $\sigma_{sep}>0$ and $N_{sep}\geq 2$ such that, for all $h\in\text{Reg}(F)$, we have, for any interval $I$ of size $|I|\leq\sigma_{sep}$:
$\forall n\geq N,\forall y\in F,\ \mathcal{L}_{\phi}^{n}(h\mathbf{1}_{I})(y)\leq\frac{h(y)}{4C_{\phi}}.$
Proof.
Since $\mu$ has no atoms and is compactly supported, its cumulative distribution function is uniformly continuous. It follows that there exists a function $\varepsilon:[0,\infty)\rightarrow[0,\infty)$ with $\varepsilon(0)=0$, continuous at zero, such that $\mu(I)\leq\varepsilon(|I|)$. Define $\sigma_{sep}>0$ so small $\varepsilon(2\sigma_{sep})\leq\frac{1}{4C_{\phi}^{2}}$. Now choose $h\in\text{Reg}(F)$, $|I|\leq\sigma_{sep}$. We can then bound:
$\mathcal{L}_{\phi}^{n}(h\mathbf{1}_{I})\leq C_{\phi}h(y)\mathcal{L}_{\phi}^{n}(\mathbf{1}_{I})\leq C_{\phi}h(y)\mathcal{L}_{\phi}^{n}(\chi).$
Where $\chi$ is smooth, takes value in $[0,1]$ and is supported in an interval of size $\sigma_{sep}$. By Perron-Frobenius-Ruelle theorem [14], we find, for $n\geq N$ where $N$ is some constant that only depends on the $IFS$: $\mathcal{L}_{\phi}^{n}(\chi)\leq\frac{1}{8C_{\phi}^{2}}+\int\chi d\mu\leq\frac{1}{4C_{\phi}^{2}}$. The desired bound follows. ∎
Lemma 2.6.
Define the $N$-th exponential moment function by $\mathcal{M}_{N}(t):=\sup_{x,y}\sum_{\mathbf{\mathbf{a}}\in\Sigma^{N}}\omega_{\mathbf{a}}(x)g_{\mathbf{a}}^{\prime}(y)^{-t}$, which is finite for small enough $t>0$ by hypothesis, and satisfies $\mathcal{M}_{N}(0)=1$. Choose $s_{\mu}\in(0,1)$ such that $\mathcal{M}_{N_{sep}}(s_{\mu})\leq 2$. Define further $C_{reg}:=2C_{\phi}\sigma_{sep}^{-1}$ from Lemma 2.5. Then for all $h\in\text{Reg}(F)$, for all interval $I$, we have:
$\forall y,\ \mathcal{L}_{\phi}^{n}(h\mathbf{1}_{I})(y)\leq h(y)\cdot C_{reg}\Big(|I|^{s_{\mu}}+\frac{1}{2^{n/N_{sep}}}\Big)$
Proof.
The proof is done by induction on $n\geq 0$. If $n\leq N_{sep}$, then the bound is trivial, since $\varphi$ is normalized, and since $h\in\text{Reg}(F)$: $\mathcal{L}_{\phi}^{n}(h\mathbf{1}_{I})\leq C_{\phi}h(y)\mathcal{L}_{\phi}^{n}(\mathbf{1}_{F})\leq 2C_{\phi}\Big(|I|^{s_{\mu}}+\frac{1}{2^{n/N}}\Big)$.
Suppose then that $n\geq N_{reg}+1$. If $\sigma>\sigma_{sep}/2$, the bound is also trivial. So suppose that $\sigma<\sigma_{sep}/2$.
We have:
$\mathcal{L}_{\phi}^{n}(h\mathbf{1}_{I})(y)=\sum_{\mathbf{c}\in\Sigma^{n}}\omega_{\mathbf{c}}(y)h(g_{\mathbf{c}}y)\mathbf{1}_{I}(g_{\mathbf{c}}(y))$
$\sum_{\mathbf{a}\in\Sigma^{n-N}}\sum_{\mathbf{b}\in\Sigma^{N}}\omega_{\mathbf{b}}(g_{\mathbf{a}}x)\omega_{\mathbf{a}}(x)h(g_{\mathbf{b}\mathbf{a}}y)\mathbf{1}_{I}(g_{\mathbf{b}\mathbf{a}}y).$
Now notice that $\mathbf{1}_{I}(g_{\mathbf{b}\mathbf{a}}y)=\mathbf{1}_{I_{\mathbf{b}}}(g_{\mathbf{a}}y)\mathbf{1}_{\tilde{I}}(g_{\mathbf{b}}(y))$ where $\tilde{I}$ is an interval of size $|\tilde{I}|<\sigma_{reg}$, since $N_{reg}$ is large, and where $I_{\mathbf{b}}$ is an interval of size $|g_{\mathbf{b}}^{\prime}|^{-1}|I|$. In particular:
$\mathcal{L}_{\phi}^{n}(h\mathbf{1}_{I})(y)=\sum_{\mathbf{b}\in\Sigma^{N}}\sum_{\mathbf{a}\in\Sigma^{n-N}}\omega_{\mathbf{a}}(y)\omega_{\mathbf{b}}(g_{\mathbf{a}}(y))h(g_{\mathbf{b}}g_{\mathbf{a}}(y))\mathbf{1}_{I_{\mathbf{b}}}(g_{\mathbf{a}}(y))\mathbf{1}_{\tilde{I}}(g_{\mathbf{b}}(y))$
$\sum_{\mathbf{b}\in\Sigma^{N}}\mathbf{1}_{\tilde{I}}(g_{\mathbf{b}}(y))\mathcal{L}^{n-N}_{\phi}(h_{\mathbf{b}}\mathbf{1}_{I_{\mathbf{b}}})(y)$
$\leq\sum_{\mathbf{b}\in\Sigma^{N}}\mathbf{1}_{\tilde{I}}(g_{\mathbf{b}}(y))\omega_{\mathbf{b}}(y)h(g_{\mathbf{b}}y)C_{reg}\Big(|I_{\mathbf{b}}|^{s_{\mu}}+2^{-(n-N)/N}\Big)$
$\leq C_{reg}\Big(\sum_{\mathbf{b}}\omega_{\mathbf{b}}h\circ g_{\mathbf{b}}\mathbf{1}_{\tilde{I}}\circ g_{\mathbf{b}}\Big)^{1/2}\Big(\sum_{\mathbf{b}}\omega_{\mathbf{b}}h\circ g_{\mathbf{b}}|g_{\mathbf{b}}^{\prime}|^{-2s_{\mu}}\Big)^{1/2}|I|^{s_{\mu}}+C_{reg}h(y)2^{-n/N}$
$\leq C_{reg}\Big(\frac{h(y)^{1/2}}{2C_{\phi}^{1/2}}(C_{\phi}h(y))^{1/2}M_{N_{sep}}(2s_{\mu})|I|^{s_{\mu}}+h(y)2^{-n/N}\Big)$
$\leq h(y)\cdot C_{reg}\Big(|I|^{s_{\mu}}+2^{-n/N}\Big).$
We can now prove Proposition 2.2 using Lemma 2.6.
Proof of Proposition $2.$
By Lemma 2.6, we have, for all $n$:
$\mu(I)=\int_{F}\mathbf{1}_{I}d\mu=\int_{F}\mathcal{L}_{\varphi}^{n}(\mathbf{1}_{I})d\mu\leq C_{reg}\Big(|I|^{s_{\mu}}+2^{-n/N_{sep}}\Big).$
Taking the limit in $n\rightarrow\infty$ gives $\mu(I)\leq C_{reg}|I|^{s_{\mu}}$. ∎